OpenStudy (anonymous):

A monument is built out of toothpicks. There are 30 rows and each row has seven more toothpicks than the row above. The top row has five toothpicks. How many toothpicks are on the bottom row? Can someone explain how I would solve this I don't understand

OpenStudy (anonymous):

You could write a function for the amount of toothpicks per row: Row 1 (the top) has 5 and you add 7 each time: t = 7(r) - 2 where t = # of toothpicks and r = row

OpenStudy (lilg132):

why is it - 2?

OpenStudy (anonymous):

To make row 1 have 5 toothpicks

OpenStudy (lilg132):

so the answer is 208 toothpicks?

OpenStudy (lilg132):

t = 7(30) - 2

OpenStudy (anonymous):

How did you get the -2 though

OpenStudy (anonymous):

xjai, did you give up on the other geometric series problem I was helping you with?

OpenStudy (anonymous):

That is the answer 208

OpenStudy (anonymous):

The 7 is the amount of toothpicks added each time and the -2 is because the first row has 5 toothpicks in it (5 - 7 = -2)

OpenStudy (lilg132):

no the answer is t = 5 + (r - 1)d t = 5 + (30 - 1)7

OpenStudy (lilg132):

well the answer is the same sorry but easier equation to work with and to understand with

OpenStudy (anonymous):

To-MA-to, to-MAY-to

OpenStudy (lilg132):

where t = total number of toothpicks on last row r = the amount of rows d = how much more in each row

OpenStudy (anonymous):

Ok, I was lost for a minute but the equation Lil gave made me understand it better